OpenStudy (anonymous):

What is the quadratic function that is created with roots -3 and 1 and a vertex at (-1, -8)?

OpenStudy (anonymous):

f(x)=2(x+3)(x-1)

OpenStudy (anonymous):

f(x) = -2x2 + 4x − 6 f(x) = 2x2 + 4x − 6 f(x) = 2x2 + 4x + 6 f(x) = -2x2 − 4x − 6

OpenStudy (anonymous):

those are the only answers i can choose

OpenStudy (nurali):

y = a(x + 3)(x - 1) -8 = a(-1 + 3)(-1 - 1) -8 = -4a 2 = a y = 2(x + 3)(x - 1) y=2(x^2+2x-3) y=2x^2+4x-6